# Harmony in 10-TET: From Parallel Universes to the Desargues Configuration

### v1

@@ 7.3. Algebraic Characterization and Graph Invariants
7.3.1. The Grand Tour: A Hamiltonian Cycle
A specific Hamiltonian cycle can be generated by the operation sequence $(R\cdot L)^{10}$. Starting from $D_{0}$, this sequence visits every chord exactly once before returning to the start. The progression effectively transposes the chord by a minor third (interval class 3) at each $L$-step:
$D_{0}\xrightarrow{R}M_{7}\xrightarrow{L}D_{3}\xrightarrow{R}M_{0}\xrightarrow{L}D_{6}\xrightarrow{R}M_{3}\dots\xrightarrow{L}D_{0}$
Geometrically, this path zig-zags between the “Blue” and “Red” universes, weaving them into a single continuous thread.
$D_{9}$$M_{3}$$D_{3}$$M_{7}$$D_{7}$$M_{1}$$D_{1}$$M_{5}$$D_{5}$$M_{9}$$D_{6}$$M_{0}$$D_{0}$$M_{4}$$D_{4}$$M_{8}$$D_{8}$$M_{2}$$D_{2}$$M_{6}$$D_{0}$(a) Concentric Visualization
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$D_{0}$(b) Circular Visualization
Figure 12. A Hamiltonian cycle ("The Grand Tour") visualized in two distinct layouts. The sequence $(RL)^{10}$ visits every chord exactly once. The path alternates between Identity Wormholes (dashed gold) and Leading-Tone steps (solid orange). (a) In the concentric layout, arrows indicate traversal order. (b) In the standard circular layout, the path forms a complex star-polygon.
7.4. The Classification Problem: Kantor, Martinetti, and the Missing Graph
Having established that our graph is a connected $(10_{3})$ configuration (in the combinatorial sense of being 3-regular with 10 “points” and 10 “lines”), we must check it against the historical classification of such structures.
In the late 19th century, Kantor (1881) [9] and Martinetti (1887) [10] rigorously classified the ten non-isomorphic connected $(10_{3})$ configurations. A comparison of invariants, however, yields a discrepancy: our functional 10-TET graph appears nowhere on this list.
This absence is not due to an error in the graph construction, nor an incompleteness in the classical list, but rather a divergence in the axiomatic definition of a “configuration.” The Kantor-Martinetti classification is grounded in projective geometry, which imposes the Axiom of Incidence:
Two distinct points are incident with at most one common line.
In graph-theoretical terms, this axiom forbids the existence of “digons” or “squares” (cycles of length 4). If a graph contains a 4-cycle ($P_{1}-L_{1}-P_{2}-L_{2}-P_{1}$), it implies that lines $L_{1}$ and $L_{2}$ intersect twice (at $P_{1}$ and $P_{2}$), which is geometrically degenerate. Therefore, all ten classical configurations possess a girth $g\geq 6$.
Since our graph possesses a girth $g=4$ (as shown in Section 7.3), it violates the geometric axiom of incidence. It is, therefore, a “degenerate” configuration in the eyes of classical geometry, but a perfectly valid (and physically meaningful) object in topological graph theory.
7.5. Identification: The Generalized Petersen Graph
To identify the precise isomorphism class of our graph, we turn to the generalized classification of cubic graphs provided by the Petersen $GP(10,k)$ family [2, 5].
Theorem 4.
The functional 10-TET graph is isomorphic to the Decagonal Prism Graph, also classified as the Generalized Petersen Graph $GP(10,1)$.
Proof.
The Generalized Petersen Graph $GP(n,k)$ is defined by vertices $\{u_{0},\dots,u_{n-1}\}\cup\{v_{0},\dots,v_{n-1}\}$ and edges:
$(u_{i},u_{i+1}),\quad(u_{i},v_{i}),\quad(v_{i},v_{i+k})$
By inspection of the 10-TET graph properties:
  1. (1)
    We identify two disjoint 10-cycles formed by the $P$ and $L$ operations: the Blue Ring ($u_{i}$) and the Red Ring ($v_{i}$).
  2. (2)
    The Wormhole operation $R$ connects each node in the Blue Ring to a unique node in the Red Ring, establishing the “spokes” $(u_{i},v_{i})$.
  3. (3)
    Unlike the Desargues graph (which is $GP(10,3)$), our rings preserve local connectivity without skipping steps in the permutation cycle.
Thus, setting $n=10$ and $k=1$, we observe that the structure is a prism with bases connected by identity wormholes. This confirms the isomorphism to $GP(10,1)$. ∎
8. The “Wide” Choice of Thirds: A Geometric Revelation
We now return to the singular case identified in Section 5.4: the “Wide” 10-TET system defined by the parameters $(t,s)=(6,1)$. While the earlier sections focused on the topological continuity of the “narrow” systems (where $\Delta=1$), the $\Delta=5$ case offers a different kind of beauty. It is not merely another valid tuning; as we shall see, it turns out to be the only configuration in the 10-TET universe that realizes a fully vertex-transitive Levi graph structure.
In the following analysis, we demonstrate that this musical system is not just a graph of chords, but an exact musical isomorphism of the famous Desargues configuration $(10_{3})_{1}$. This result bridges the gap between harmonic theory and projective geometry, revealing that the most symmetric musical choice in 10-TET corresponds to one of the most fundamental theorems in mathematics.
@@ 8.4. The Levi Graph
8.5. Topology of Cycles: Rigorous Proof of Girth 6
@@ 8.6. Parity Constraint
8.7. The "Shift Method" for Even Cycles
@@
    PL (via $M_{r}$): $D_{r}\to M_{r}\to D_{r-6}$. Shift: $\mathbf{+4}$ (since $6\equiv 4$).
@@
    RL (via $M_{r-1}$): $D_{r}\to M_{r-1}\to D_{r-7}$. Shift: $\mathbf{+3}$ (inverse of LR).
8.8. Proof that Length 4 is Impossible
@@
  • $4+6=10$. This looks promising! However, shift +4+4 is $PL$ and shift +6+6 is $LP$. The sequence would be $PL$ followed by $LP$, i.e., $P\to L\to L\to P$. Since $L^{2}=Identity$, this collapses to $P\to P=Identity$. This is a backtracking path, not a cycle.
  • The same logic applies to $1+9=10$ (PR and RP) and $3+7=10$ (RL and LR).
@@
8.9. Proof that Length 6 Exists
A cycle of length 6 requires **three** double-steps. We look for three shifts that sum to 10. We find the solution:
$1+6+3=10$
Mapping these numbers back to operations:
  • Shift $\mathbf{+1}$: Operation $PR$.
  • Shift $\mathbf{+6}$: Operation $LP$.
  • Shift $\mathbf{+3}$: Operation $RL$.
Linking these sequentially creates the path: $(PR)\to(LP)\to(RL)$. This corresponds to the sequence $P,R,L,P,R,L$. This is exactly the $(PRL)^{2}$ cycle.
8.10. Chirality
Because the operations do not commute, the order matters.
  • **Right-Handed Cycle $(PRL)^{2}$:** Uses the sum $1+6+3$.
  • **Left-Handed Cycle $(PLR)^{2}$:** Corresponds to the shifts $4+9+7=20\equiv 0$.
These are the two shortest paths through the harmonic space, visualized below.
@@ 8.10. Chirality
8.11. Geometric Identification: The Desargues Configuration
The structural properties of the derived Tonnetz lead to a significant geometric identification. We have established that the graph is:
@@ 8.11. Geometric Identification: The Desargues Configuration
    **Cubic (3-valent):** Every chord has exactly 3 neighbors.
@@
    **Bipartite:** Connections only exist between Major and Minor chords.
@@
    **Order 20:** There are 10 Major and 10 Minor chords.
@@
    **Girth 6:** The shortest cycle has length 6.
@@
    **Vertex-Transitive:** Due to the cyclic generation of chords via $\mathbb{Z}_{10}$ and the uniform definition of transformations $P,L,R$, the graph exhibits a high degree of symmetry. Mapping $D_{r}\to D_{r+k}$ implies a corresponding automorphism of the entire graph.
In the classification of combinatorial configurations, a system of $n$ points and $n$ lines where each point lies on 3 lines and each line contains 3 points is denoted as a $(n_{3})$ configuration. The Levi graph of such a configuration connects points to lines, resulting in a bipartite cubic graph of $2n$ vertices.
For $n=10$, there are exactly 10 distinct combinatorial types of $(10_{3})$ configurations (see [1] Table 1, p.333, or [14], Section 5.3.4, p. 176 - 179). Among these, **only one** possesses a vertex-transitive Levi graph: the **Desargues Configuration**, usually denoted as $(10_{3})_{1}$.
@@
The Levi graph of the Wide 10-TET system ($\Delta=5$) is isomorphic to the **Desargues Graph**.
Proof.
The Desargues graph is the unique symmetric cubic graph on 20 vertices. Since our harmonic graph is generated by a vertex-transitive group action on $\mathbb{Z}_{10}$ and matches the order and valency requirements, it must be isomorphic to the Desargues graph. ∎
This identification is profound. The Desargues configuration is the geometric embodiment of Desargues’ Theorem, a fundamental result in projective geometry. Unlike other configurations that may require specific fields, the Desargues configuration can be realized in any Desarguesian projective plane $PG(2,\mathbb{K})$, including the real projective plane and finite fields $\mathbb{K}$ such as $\mathbb{F}_{5}$.
Thus, the "Wide" 10-TET harmonic system is not merely a graph of chords; it is a musical realization of one of the most classic structures in projective geometry. The Major chords can be viewed as "Points" and the Minor chords as "Lines" (or vice versa) in a projective configuration where harmonic relations define incidence.
@@ 9. The Tritone Solution: System (5,2)
In Section 6.3, we briefly encountered the “neutral” solution to the Pythagorean constraints, characterized by the difference $\Delta=3$.
@@
This yields the solution $t=5$ and $s=2$. Although we initially set this aside due to the acoustic strangeness of treating the tritone ($t=5$) as a consonance, the topological robustness of this system merits a full analysis. As we shall see, this system serves as a fascinating “middle ground” between the disconnected Acoustic system ($\Delta=1$) and the highly symmetric Wide system ($\Delta=5$).
@@ 9.5. The Tritone Tour
$(P\cdot R)^{4}\cdot L\cdot(R\cdot P)^{4}\cdot R\cdot L\cdot R$
@@
    Bridge: $D_{8}\xrightarrow{L}M_{3}$ (Switch to Odd).
@@
    Return: $M_{1}\xrightarrow{R}D_{3}\xrightarrow{L}M_{8}\xrightarrow{R}D_{0}$.
@@
9.6. Identification with the Generalized Petersen Graph
Having analyzed the local and global properties of the $(5,2)$ system, we can now formally identify its underlying graph structure.

### v2

@@ 7.3. Algebraic Characterization and Graph Invariants
7.3.1. Hamiltonian Connectivity: Two Paths
The graph is Hamiltonian, meaning there exists a path that visits every chord exactly once before returning to the start. We identify two distinct Hamiltonian cycles generated by alternating operations, each revealing a different aspect of the system’s symmetry.
The Perimeter Cycle ($R\cdot P$).
The first cycle corresponds to the perimeter arrangement shown previously in Figure 14(b). It is generated by alternating the Parallel ($P$) and Wormhole ($R$) transformations:
$C_{PR}=(R\cdot P)^{10}$
Starting from $D_{0}$, the sequence proceeds: $D_{0}\xrightarrow{P}M_{0}\xrightarrow{R}D_{3}\xrightarrow{P}M_{3}\xrightarrow{R}D_{6}\dots$. This path traces the local inversion ($P$) followed by a global shift ($R$), naturally unfolding the graph’s maximal symmetry into a ring.
The Grand Tour ($L\cdot R$).
We can construct a second, distinct Hamiltonian cycle by alternating the Wormhole ($R$) and Leading-Tone ($L$) operations:
$C_{LR}=(L\cdot R)^{10}$
This path effectively transposes the chord by a minor third (interval class 3) at each step:
$D_{0}\xrightarrow{R}M_{7}\xrightarrow{L}D_{3}\xrightarrow{R}M_{0}\xrightarrow{L}D_{6}\dots$
While the algebraic formula is as simple as the perimeter cycle, its geometric realization in the standard sequential layout ($D_{0},M_{0},D_{1}\dots$) is strikingly different. Instead of stepping between neighbors, the Grand Tour weaves a complex, star-like pattern, jumping across the circle to stitch the entire manifold together.
$D_{9}$$M_{3}$$D_{3}$$M_{7}$$D_{7}$$M_{1}$$D_{1}$$M_{5}$$D_{5}$$M_{9}$$D_{6}$$M_{0}$$D_{0}$$M_{4}$$D_{4}$$M_{8}$$D_{8}$$M_{2}$$D_{2}$$M_{6}$$D_{0}$(a) Concentric Flow
$D_{0}$$M_{0}$$D_{1}$$M_{1}$$D_{2}$$M_{2}$$D_{3}$$M_{3}$$D_{4}$$M_{4}$$D_{5}$$M_{5}$$D_{6}$$M_{6}$$D_{7}$$M_{7}$$D_{8}$$M_{8}$$D_{9}$$M_{9}$$D_{0}$(b) Standard Layout
$D_{0}$$M_{7}$$D_{3}$$M_{0}$$D_{6}$$M_{3}$$D_{9}$$M_{6}$$D_{2}$$M_{9}$$D_{5}$$M_{2}$$D_{8}$$M_{5}$$D_{1}$$M_{8}$$D_{4}$$M_{1}$$D_{7}$$M_{4}$$D_{0}$(c) Perimeter Layout
Figure 16. The "Grand Tour" Hamiltonian Cycle $(L\cdot R)^{10}$. (a) Concentric flow. (b) Star-polygon structure in standard layout. (c) Perimeter layout: the nodes are re-indexed such that the $(L\cdot R)^{10}$ Hamiltonian cycle forms the outer boundary. The remaining $P$ operations are shown as internal grey chords to represent the full Levi graph structure for completeness.
8. The “Wide” Choice of Thirds: A close Desargues cousin
We now return to the singular case identified in Section 5.4: the “Wide” 10-TET system defined by the parameters $(t,s)=(6,1)$. While earlier sections focused on the topological continuity of the “narrow” systems (where $\Delta=1$), the $\Delta=5$ case offers a different kind of beauty. It satisfies the rigorous conditions of a linear incidence geometry, turning out to be one of the $10_{3}$ configurations—structures first encountered in the 17th century in the Desargues Theorem of projective geometry. At this point, a historical remark is in order.
Remark 1.
The Evolution of $10_{3}$ Configurations:
The concept of a $10_{3}$ configuration first appeared in the 17th century in the work of Girard Desargues [7]. The incidence structure of the 10 points and 10 lines involved in his theorem was recognized to possess a remarkable self-dual property, allowing the roles of points and lines to be interchanged.
For more than two centuries following its discovery, the Desargues configuration remained the unique prototype of such a system. The abstract combinatorial definition of a $10_{3}$ configuration was formulated much later. In 1881, Seligmann Kantor [13] broke this solitude by discovering a second, non-isomorphic $10_{3}$ configuration. However, Kantor’s configuration (often denoted $(10_{3})_{8}$ in older literature) posed a geometric anomaly: unlike the Desargues configuration, it cannot be realized by points and straight lines in the real projective plane.
Kantor’s discovery spurred a century-long effort to classify all possible $10_{3}$ structures. Over time, other non-equivalent and realizable configurations were identified. The classification was finally completed and proven in the late 20th century. In 2000, Betten, Brinkmann, and Pisanski published the definitive catalogue [3], proving there are precisely 10 distinct isomorphism classes.
In this classification, the configurations are distinguished, in particular, by their automorphism groups. Remarkably, the Desargues configuration is identified as type $(10_{3})_{8}$, possessing the largest group of symmetries (flag-transitive, but not cyclic). The only $10_{3}$ configuration in the entire classification that admits a cyclic symmetry group acting transitively on the points is designated as:
Configuration $(10_{3})_{3}$
This implies that any “Harmonic” realization generated purely by a cyclic group $C_{10}$ is combinatorially isomorphic to configuration $(10_{3})_{3}$, a cyclic cousin of the classical Desargues configuration $(10_{3})_{8}$.
Closing this historical aside, it is worthwhile to note that although this configuration is only a relative of the Desargues configuration, it can also be realized as a system of 10 straight lines and 10 points on the real plane. It obeys the rules that (a) each point meets three lines, (b) each line contains three points, and (c) the axioms of linear incidence geometry are satisfied (specifically, that any two distinct points determine at most one line).
In the following analysis, we demonstrate that the “Wide” musical system constitutes a valid combinatorial configuration of type $(10_{3})_{3}$. This result bridges the gap between harmonic theory and combinatorial geometry, revealing that this specific choice of intervals generates a graph with a girth of $g=6$, satisfying the Axiom of Incidence required for a true point-line geometry.
@@ 8.4. The Levi Graph
8.5. Topologgy of Cycles: Girth 6
@@ 8.5.1. Parity Constraint
8.5.2. The "Shift Method" for Even Cycles
@@
    PL: $D_{r}\to M_{r-1}\to D_{r-1}$. Shift: $\mathbf{9}$ (since $1\equiv 9$).
@@
    RL: $D_{r}\to M_{r-1}\to D_{r+3}$. Shift: $\mathbf{3}$.
8.5.3. Proof that Length 4 is Impossible
@@
  • $4+6=10$. This looks promising! However, shift $4$ is $RP$ and shift $6$ is $PR$. The sequence would be $PR$ followed by $RP$, i.e., $RP\cdot PR=RP^{2}R$. Since $P^{2}=Identity$, this collapses to a $P\to P$ path. This is a backtracking path, not a cycle. The same story is when we do $RP$ followed by $PR$, since $R^{2}=Identity$.
  • The same logic applies to $1+9=10$ (PL and LP) and $3+7=10$ (RL and LR).
@@
8.5.4. Proof that Length 6 Exists
It is enough to show at least one such cycle. Examples of two paths of length 6 of oposite chirality through the harmonic space is visualized below.
@@ 8.5.4. Proof that Length 6 Exists
8.6. Identification as a Cyclic Configuration
We now list structural properties of the derived Tonnetz which are easy to detect. We first determine the “offsets” – the specific intervals that connect Major chords to Minor chords in the graph.
By examining the common tones defined in Section 8.3, we observe that every Major chord $D_{r}$ shares two notes with exactly three Minor chords:
@@ 8.6. Identification as a Cyclic Configuration
    $M_{r+\textbf{0}}$ (sharing $\{r,r+7\}$): An offset of 0.
@@
    $M_{r+\textbf{6}}$ (sharing $\{r+6,r+7\}$): An offset of 6.
@@
    $M_{r+\textbf{9}}$ (sharing $\{r,r+6\}$): An offset of 9.
Thus, the connection set for the system is $S=\{0,6,9\}$. This set of jumps defines the graph’s invariants:
@@
    Cubic (3-valent): Every chord has exactly 3 neighbors determined by the offsets.
@@
    Bipartite: Connections only exist between Major and Minor chords.
  3. (3)
    Order 20: There are 10 Major and 10 Minor chords.
  4. (4)
    Girth 6: The shortest cycle has length 6. A concrete example of such a hexagon, generated by the alternating sum of offsets (96+09+60=09-6+0-9+6-0=0), is:
    $D_{0}\xrightarrow{9}M_{9}\xrightarrow{6}D_{3}\xrightarrow{0}M_{3}\xrightarrow{9}D_{4}\xrightarrow{6}M_{0}\xrightarrow{0}D_{0}$
    (Arrows denote the offset magnitude connecting the nodes; this is the opposite 6-cycle to the violet cycle in Figure 19).
  5. (5)
    Hamiltonicity System is visibly Hamiltonian having all points on the perimeter of its Levi graph connected (seeFigure 18).
  6. (6)
    Cyclic Symmetry: The graph admits a cyclic group action $C_{10}$, that maps $D_{r}\to D_{r+k}$ and $M_{r}\to M_{r+k}$ and which is generated by a clockwise rotation by 36 degrees: $D_{r}\to D_{r+1}$ and $M_{r}\to M_{r+1}$. This acttion is transitive and preserves the offset structure. So the cyclic group $C_{10}$ is visibly a (sub)group of the full musical symmetry group of the graph acting transitively on it.
With the geometric invariants and the automorphism group of the Wide system now fully characterizated, we return to the classification context established in Remark 1.
Recall that the census by Betten, Brinkmann, and Pisanski identifies the Cyclic Configuration as a singular entity: it is the only one, denoted by $(10_{3})_{3}$, among the ten distinct classes to admit a cyclic symmetry group acting transitively on the points. Since our analysis has demonstrated that the Wide 10-TET system is both a valid $(10_{3})$ configuration and inherently generated by the cyclic group $C_{10}$, the identification follows immediately from this uniqueness.
@@
The Levi graph of the Wide 10-TET system ($\Delta=5$) is isomorphic to the **Unique Cyclic Configuration $(10_{3})**$, labeled as $(10_{3})_{3}$ in the Betten census.
This identification reveals the precise nature of the “Wide” system. Unlike the degenerate $(4,3)$ system (which contains squares and thus fails the axioms of a linear geometry), the $(6,1)$ system satisfies the condition of Girth 6, making it a valid configuration of points and lines.
It is a **Self-Dual** geometry: the graph structure is invariant under the exchange of Major and Minor chords. However, its most significant feature is its distinctness from the Desargues configuration known from the Desargues theorem in projective geometry. While Desargues is highly symmetric (Group symmetry order 240, flag-transitive), the Wide system is the **unique** configuration among the ten combinatorial types of $(10_{3})$ that admits a cyclic automorphism group. Thus, it represents the singular intersection between the axioms of combinatorial geometry and the cyclic symmetry inherent in musical theory.
@@ 9. The Tritone Solution: System (5,2)
In Section 6.3, we encountered the “neutral” solution to the Pythagorean constraints, characterized by the difference $\Delta=3$.
@@
This yields the solution $t=5$ and $s=2$.
Mathematically, this system is characterized by its extreme symmetry. Analysis of the underlying graph reveals a Girth of 4, indicating the immediate presence of squares (4-cycles) in the chordal network.
@@ 9.5. The Tritone Tour
${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}L}\cdot R\cdot(P\cdot R)^{4}\cdot{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}L}\cdot R\cdot(P\cdot R)^{4}$
@@
    Bridge: $D_{8}\xrightarrow{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}L}M_{3}$ (Switch to Odd).
@@
    Return Bridge: $D_{3}\xrightarrow{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}L}M_{8}$ (Switch to even).
@@
10. Identifications of the three 10-TET harmony systems
In Sections 7.3 and 9.4 we showed that the “Acoustic” and “Tritone” systems have girth 4. As such they can not be interpreted as configurations in a classical sense, since in the interpretation of major and minor chords as respective points and lines, there are pair of points in them that meet more than one, line, and there lines that meet ech other in more than one point. Therefore these systems can not be found in the list of ten $10_{3}$ configurations provided by [3]. After analyzing those invariants of graphs of these systems, which we were able to determine, we believe that both of them belong to the family of Circulant Graphs [4, 8]. Roughly: A graph $Ci_{n}(s_{1},\dots,s_{k})$ is circulant if it has $n$ vertices with edges determined by a set of jumps. In the case of our two systems and their bipartite levi graphs, we map the 20 chords to a cycle $\mathbb{Z}_{20}$ such that Major chords are at even positions ($2i$) and Minor chords at odd positions ($2i+1$).
10.1. Circulant graphs
Before stating the main result for this system, we must rigorously define the target family of graphs.
Definition 1 (Circulant Graph).
[1, 4, 8] A Circulant Graph $Ci_{n}(s_{1},\dots,s_{k})$ is a graph with $n$ vertices labeled $0,1,\dots,n-1$, where two vertices $u$ and $v$ are adjacent if and only if:
$u-v\equiv\pm s_{j}\pmod{n}$
for some $s_{j}$ in the generating set $S=\{s_{1},\dots,s_{k}\}$.
10.1.1. The “Acoustic” system $(4,3)$